Ram will be 5 times as old as he is now after 8 years .how many years hence will he be 10 times as old as he is now.
Answer by ankor@dixie-net.com(22740)

Let r = his present age
:
Ram will be 5 times as old as he is now after 8 years
5r = r + 8
5r - r = 8
4r = 8
r = 2 yrs now
:
.how many years hence will he be 10 times as old as he is now.
let y = no. of yrs for this to be true
10(2) = (y+2)
20 = y + 2
20 - 2 = y
y = 18 yrs this will be true